OpenStudy (anonymous):

What is the midpoint of (–4, 4) and (2, –4)? (0, 0) (–6, 8) (–1, 0) (–2, 0)

OpenStudy (anonymous):

graph the point and you can see for your self. ;)

OpenStudy (anonymous):

\[ x=\frac{x_1+x_2}2\\ y=\frac{y_1+y_2}2\\ \]

OpenStudy (anonymous):

(–4, 4) and (2, –4)? \[ x=\frac{-4+2}{2}=-1\\ y=\frac{-4+4}{2}=0 \] The answer is (-1,0)

OpenStudy (rayford):

very good! :[]